A mixture containing 2.17 g each of CH4(g), C2H4(g) and C4H10(g) is contained in a 1.50 L flask at a temperature of 35°C. (a) Calculate the partial pressure of each of the gases in the mixture.

PCH4 = ? atm

PC2H4 = ? atm

PC4H10 = ? atm

(b) Calculate the total pressure of the mixture. ? atm
Answer #1

moles CH4 = 2.17 g = 0.136 - 16g/mol C₂H4 = 2.17 g 0.0775 28g1mo Cu Hior 2.179 co.0388 - 56g/mol n total = 0.2523 mole fracti